OpenStudy (anonymous):

Hi. I trying to work out the integral of 1/(sqrt(2x-x^2)), but I'm moving too slowly. Can anyone give me any tips on how to move faster doing questions like this? @MIT 18.02 Multiva…

OpenStudy (jamesj):

Write 2x - x^2 = 1 - (x-1)^2 Now substitute x - 1 = sin t

OpenStudy (anonymous):

this looks like a setup for competing the square. since you are close with \[2x-x^2\] i would change it to ... what jamesj said

OpenStudy (anonymous):

but i would not make a trig sub if you remember that \[\frac{1}{\sqrt{1-x^2}}=\frac{d}{dx}\sin^{-1}(x)\] you are more than half way home

OpenStudy (jamesj):

If you do that you should see that \[\frac{1}{\sqrt{2x - x^2}} = \frac{1}{\sqrt{1 - \sin^2 t}} = \frac{1}{\cos t}\]

OpenStudy (jamesj):

yet if you don't know that result this is how you prove it.

myininaya (myininaya):

satellite likes to take the easy way out all the time

OpenStudy (jamesj):

(or rather, this is one way to prove it.)

OpenStudy (anonymous):

it never hurts to remember a few things. like the derivative of inverse trig functions (as least arcsine and arcosine) and derivative of roots and reciprocals. it is like knowing that 7 times 8 is 56 or whatever it is

myininaya (myininaya):

i like making my students proving it each time i figure they learn more that way and also their algebra sucks so more algebra practice means they will get better at it right?

OpenStudy (anonymous):

you are mean

myininaya (myininaya):

formulas cheat them

OpenStudy (anonymous):

you should make them prove each time that \[\frac{d}{dx}x^n=nx^{n-1}\] because that is the only thing they are going to remember after 2 years trust me

myininaya (myininaya):

lol i tried to make them prove the product rule but thats hard for them

OpenStudy (anonymous):

and then you should make them prove that \[(f\circ g)'=f'(g)g'\] (not to bring up a sore subject...)

OpenStudy (jamesj):

I actually this there's an important difference between the derivative of x^n and arcsin x. Which is how to deal with constants. We all know what to do with the constants when integrating or differentiating cx^n But unless you've done it recently, it's hard to remember exactly what the result is when you integrate \[\frac{1}{\sqrt{a^2 - x^2}} \ \ or \frac{1}{\sqrt{1-a^2x^2}}\]

OpenStudy (anonymous):

i will agree with that because unless i have done them recently i can integrate literally nothing